OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} \frac{n^{2}+1}{n^{3} +1}\] the way I would solve it is \[\frac{n^{2}/n^{2}+1/n^{2}}{n^{3}/n^{2} +1/n^{2}}\] \[\frac{1+1/n^{2}}{n+1/n^{2}}\] and as \[ n\rightarrow \infty\] we get \[1/\infty \] =0 put the answer is: divergent, so I guess my calculations are wrong

OpenStudy (anonymous):

Can you take n^3 common from both numerator and denominator???

OpenStudy (anonymous):

you mean in stead of n^3?

OpenStudy (anonymous):

i mean n^2

OpenStudy (anonymous):

Yes please...

OpenStudy (anonymous):

so \[\frac{1/n+1/n^{3}}{1+1/n^{3}}\] that would make it \[\infty/1\] divergent i guess Thanks

OpenStudy (anonymous):

1/infinity is how much Mathsofiya??

OpenStudy (anonymous):

1/infty is zero but infinity/1 = infty

OpenStudy (anonymous):

So, I think answer will be : 0/1 = 0

OpenStudy (anonymous):

what about \[a_{n}/b_{n}\] ? my book multiplied n/1 to \[a_{n}\]

OpenStudy (anonymous):

I did not get what you said... explain it more..

OpenStudy (anonymous):

The book calls it the limit comparison test \[\lim_{n\rightarrow\infty} \frac{a_{n}}{b_{n}}\] and somehow they took an n/1 and multiplied it to \[ \frac{n^{2}+1}{n^{3} +1}\]

OpenStudy (anonymous):

Well, I have no knowledge about this. Sorry, but I will search this and make you clear this one....

OpenStudy (anonymous):

the answer should be 0, you did it right o-o...